A 2kg block at rest is acted upon by a constant force of 20N for 4 seconds. Find: (a). what momentum is required by the block ? (b). Find the final velocity of the block. Show complete solutions.


1
Expert's answer
2022-01-12T08:31:28-0500

a) The required momentum can be found from the second Newton's law:


"p = F\Delta t"

where "F = 20N, \Delta t=4s". Thus, obtain:


"p = 20N\cdot 4s = 80 kg\cdot m/s"

b) The velocity can be found from the definition of the momentum:


"v = \dfrac{p}{m}"

where "m = 2kg". Thus, obtain:


"v = \dfrac{80kg\cdot m/s}{2kg} = 40m/s"

Answer. a) 80kg*m/s, b) 40 m/s.